OpenStudy (jamaljunior):

Write the equation of the line in Standard Form that passes through the points (3, 2) and (5, 8).

OpenStudy (jamaljunior):

@pooja195

563blackghost (563blackghost):

You would find the slope and y-intercept.... \(\huge{\frac{8-2}{5-3}=slope}\) What is the slope?

OpenStudy (jamaljunior):

3

OpenStudy (jamaljunior):

@563blackghost

OpenStudy (jamaljunior):

@jackthegreatest

OpenStudy (jackthegreatest):

yes slope = 3

OpenStudy (jackthegreatest):

y = mx +b 8 = 3 (5) + b solve for b

OpenStudy (mathmale):

And then, once you have 'b,' write the general equation in the form y = mx + b. Use the m and b values you've just found.

OpenStudy (jackthegreatest):

yep

OpenStudy (jamaljunior):

so 8 = 15?

OpenStudy (jackthegreatest):

8 = 15 + b

OpenStudy (jackthegreatest):

b = -7

OpenStudy (jamaljunior):

is the answer 3x + y = -7

OpenStudy (mathmale):

"8 = 15?" No, never.

OpenStudy (mathmale):

Choose either of the given points. Subst. the x- and y-coord. of this point into your own equation. Is the equation then true?

OpenStudy (jamaljunior):

isthis the answer though?

OpenStudy (jamaljunior):

3x + y = -7

OpenStudy (mathmale):

I encourage you to learn ways in which you can check your own work. For this particular case I gave you the following advice earlier: 'Subst. the x- and y-coord. of this point into your own equation. Is the equation then true?" Please follow through. subst. 5 for x and 8 for y. Is the resulting equation true or not?